# Efficient parallel scheduling for sparse triangular solvers

### v1

@@ Abstract.
We develop and analyze new scheduling algorithms for solving sparse triangular linear systems (SpTRSV) in parallel. Our approach, which we call barrier list scheduling, produces highly efficient synchronous schedules for the forward- and backward-substitution algorithm. Compared to state-of-the-art baselines HDagg [ZCL+22] and SpMP [PSSD14], we achieve a 3.24×3.24$\times$3.24 × and 1.45×1.45$\times$1.45 × geometric-mean speed-up, respectively. We achieve this by obtaining an up to $\times1$ geometric-mean reduction in the number of synchronization barriers over HDagg, whilst maintaining a balanced workload, and by applying a matrix reordering step for locality. We show that our improvements are consistent across a variety of input matrices and hardware architectures.
@@
22footnotetext: Authors are listed in alphabetical order.
@@ 1. Introduction
Systems of linear equations are ubiquitous and solving them fast numerically with high accuracy is essential to engineering, big data analytics, artificial intelligence, and various scientific subjects. Key techniques in scaling to ever larger linear systems have been exploiting the sparsity of non-zero coefficients in modern algorithms, as well as leveraging the multi-core or multi-processor architectures of high-performance computing systems. However, whilst sparsity reduces computational load, the typically irregular distribution of non-zero elements complicates the development of efficient parallel algorithms, as the lack of structure hinders workload balancing and limits the ability to minimize communication between processors.
In this paper, we concern ourselves with solving sparse triangular systems of linear equations (SpTRSV) using parallel machines; i.e., solving a linear system $Lx=b$, where $L$ is a sparse triangular matrix and $b$ is a dense vector. Although solving sparse triangular linear systems mark a special case, it often arises as an important step in procedures solving more general linear systems. Some concrete examples are (sparse) LU, QR, and Cholesky decompositions, Gauss–Seidel, and so forth. Efficient parallel-computation schedules for SpTRSV are of particular importance in applications where the same sparsity pattern is used repeatedly. Such is the case in simulations of various physical systems, for instance, ones that are based on the finite element method on a fixed mesh.
@@
Satisfying both of these needs simultaneously has proven to be challenging due to the irregular interdependence of computed values and the fine-grained nature of the problem. Early algorithms include so-called wavefront schedulers [AS89, Sal90], which repeatedly schedule all computations whose prerequisites are met, known as the wavefronts, cf. Figure 1.1(b), followed by a synchronization barrier. They, however, suffer from large overhead stemming from frequent global synchronization [PSSD14]. Similarly, early asynchronous approaches such as self-scheduling [SMB88] had the drawback of incurring overheads due to numerous fine-grained synchronizations [RG92].
@@ 1.1. Our contribution
Our work continues along the same path of reducing the number of synchronization barriers. We put forth barrier list schedulers as an effective synchronous scheduling algorithm archetype for solving sparse triangular linear systems. In particular, we present and analyze two specific algorithms that are both tailored towards the SpTRSV application. In our experiments, we establish that these algorithms produce significantly superior parallel schedules compared to the baseline methods. Specifically, our Funnel Locking algorithm achieves a reduction in execution time of 1.45×1.45$\times$1.45 × compared to SpMP and of 3.24×3.24$\times$3.24 × compared to HDagg, on the SuiteSparse Matrix Collection benchmark [DH11] used by previous studies, see Figure 1.2. On uniformly random matrices, our other algorithm, called Funnel $p$-ivotal path, achieves a similar speed-up of 1.62×1.62$\times$1.62 × compared to SpMP and 1.87×1.87$\times$1.87 × compared to HDagg in execution time.
Figure 1.2. Geometric mean and interquartile ranges of speed-ups over Serial of several algorithms on the SuiteSparse Matrix Collection [DH11] on an Intel x86 machine using 22 cores.
Our algorithms obtain these speed-ups by significantly reducing the number of synchronization barriers required: we report an up to $\times1$ reduction in the number of barriers relative to HDagg, whilst maintaining a good workload balance. The results also show that our schedulers provide consistent improvements over several different computing architectures and types of input matrices. The scheduling algorithms themselves also have a just above linear time complexity, making them viable tools for various applications.
@@
  • two novel algorithms for generating efficient parallel schedules for SpTRSV execution,
  • adaptations and extensions of previous coarsening and reordering techniques that enhance the schedules,
@@
  • a short theoretical proof showing that our DAG coarsening techniques preserve acyclicity, based on the new concept of cascades, and
@@
  • experiments confirming that the above schedulers achieve significant speed-ups over the SpMP and HDagg baselines, on various architectures and data sets, including orthogonal testing of the individual techniques proposed.
1.1.1. Barrier list schedulers
There is an extensive amount of prior work on parallel DAG scheduling in the literature. A recent survey classifies the most relevant algorithms into two major groups, list schedulers and cluster schedulers [WS18]. List schedulers [Gra69, ACD74, HCAL89, RVG02, MSQ03] typically assign priorities to each vertex and schedule them in a topological order according to these. Such algorithms are often more efficient when the number of cores is limited [WS18].
The idea of list schedulers has also been adapted recently to a setting with barrier synchronization [PAKY24]. We refer to this new approach as barrier list scheduling. Similarly to list scheduling, these algorithms maintain a set of vertices that are ready to be executed, i.e., all their parents have been computed. Whenever a core $p$ becomes free, they assign a ready vertex to $p$ that does not require new communication between cores, with a preference for vertices that can only be executed on the core $p$. The schedulers only introduce a new synchronization barrier when a given fraction of cores become idle.
In our paper, we present two barrier-list-scheduling algorithms for finding efficient parallel schedules for the SpTRSV kernel. We extend the prototype of the barrier-list-scheduling idea by Papp et al. [PAKY24] with various improvements: we refine the rule for inserting a synchronization barrier, we decrease the amount of compute resources wasted before the synchronization barriers, and we apply more advanced data structures for an efficient implementation. More importantly, the key to the behavior of a barrier-list-scheduling algorithm is the priority function, which selects the next vertex to be assigned to a free core whenever there are multiple options. In our two schedulers, these priority functions are tailored specifically towards the SpTRSV application:
  • the $p$-ivotal path scheduler aims to prioritize vertices with long outgoing paths and/or a high-number of successors in the DAG, whilst
  • the Locking scheduler aims to ensure that we can compute the highest possible number of vertices before having to insert a new synchronization barrier.
While both algorithms are similar, our experiments show that Locking performs better on the SuiteSparse Matrix Collection and $p$-ivotal path is superior on the (unstructured) random data sets.
@@ 1.1.2. Coarsening
We combine our scheduling algorithms with a DAG coarsening approach to obtain significantly smaller DAGs, whilst preferably retaining most of the original structure. Our schedulers can then be applied to the coarse graph. The resulting schedule is subsequently pulled back to the original graph to obtain the final schedule. Such a coarsening technique is widely applied in graph partitioning and scheduling tools [KAKS97, PSSS21] and has various advantages for our scheduler: it greatly reduces the size of the graph, improves data locality, and can also help reduce the number of synchronization steps.
In order to produce a valid scheduling problem, the coarsening needs to ensure that it preserves the acyclicity of the DAG. Methods that fulfill this property have been studied in several works before, see, for example, [CLB94, FER+13, HKU+17, ZCL+22] and references therein. In Section 4, we propose the concept of cascades to generalize the coarsening techniques utilized in [CLB94, §4] and [ZCL+22, §IV.B]. We then formally prove that coarsening techniques based on cascades always preserve acyclicity.
@@ 1.1.3. Reordering
Besides the scheduling algorithms above, we also apply a matrix reordering step to drastically improve data locality during the SpTRSV computation. Specifically, once the schedule is developed, we symmetrically permute the matrix according to the schedule, ensuring that values computed after each other on the same core are close to each other in this permuted representation. This idea has already been explored by Rothberg–Gupta in the 1990s [RG92], but it has not been applied in modern SpTRSV baselines, which instead try to make use of existing data locality when deriving a schedule.
@@ 1.2. Additional related work
An optimization method for parallel SpTRSV execution, orthogonal to what we have mentioned so far, is to break the lower triangular matrix into blocks and apply different algorithms and allocation of cores to each block individually depending on block type and sparsity pattern [AS89, May09, AYU21, YSAU20]. The blocks may be on the diagonal, which corresponds to a smaller instance of (sparse) triangular solve, or completely off the diagonal, which corresponds to a (sparse) matrix-vector multiplication. This has been particularly impactful for GPU implementations [LLH+16, LNL20]. Our scheduling algorithms would naturally fit in as a building block for such optimizations.
Finally, besides the forward-/backward-substitution algorithm, there are also other methods for solving sparse triangular systems, for example inversion. For this method, we mention the memory-optimal algorithms developed in previous works [AS93, PA92].
1.3. Acknowledgements
We would like to thank Weifeng Liu and Yves Baumann for stimulating conversations on this and surrounding topics.
@@ 2.1. Graph notation
We model our computations as a directed acyclic graph (DAG) $G=(V,E)$, which consists of a set of vertices $V$ and a set of directed edges $E\subseteq V\times V$. For any vertex $v\in V$, the sets of vertices $\{u\,|\,(u,v)\in E\}$ and $\{u\,|\,(v,u)\in E\}$ are called the parents of $v$ and the children of $v$, respectively. The in- and out-degree of $v$ are the number of parents and children of $v$, respectively. The degree of $v$, denoted by ${\rm deg}(v)$, is the sum of its in- and out-degree. If a vertex of the DAG has no parents/children, then it is called a source/sink vertex respectively. The DAG in our model is also complemented by vertex weights $\omega:V\rightarrow\mathbb{Z}_{>0}$ to indicate the compute cost of each operation.
@@ 2.2. Problem definition and notation
When solving sparse triangular systems, we are given a triangular matrix $A=(A_{i,j})_{i,j=1,\dots,n}\in\mathbb{R}^{n\times n}$, a dense vector $b=(b_{1},\dots,b_{n})^{T}\in\mathbb{R}^{n}$, and the goal is to solve the equation $Ax=b$ for the vector $x=(x_{1},\dots,x_{n})^{T}\in\mathbb{R}^{n}$. We assume that $A$ is non-singular, such that all its diagonal elements are non-zero. In case of a lower triangular matrix $A$, there is a natural forward-substitution algorithm for the problem, which iterates through the rows of $A$ in order and computes the values of $x$ as $x_{1}=\tfrac{b_{1}}{A_{1,1}}$, $x_{2}=\tfrac{b_{2}-A_{2,1}x_{1}}{A_{2,2}}$, and in general, as
@@
The case of an upper triangular matrix $A$, a backward-substitution algorithm follows symmetrically in the reverse direction.
@@
The parallel execution of this DAG then directly corresponds to a parallel execution of the SpTSRSV and many previous works found it more convenient to discuss their scheduling methods for this problem using this DAG representation.
@@
The total cost of a schedule is determined by the workload balance within each superstep and the number of synchronization barriers. The original BSP model includes also communication volume in its cost function. For the SpTRSV application, however, the communication happens in parallel to the computation and resolving the synchronizations. Hence, the latter two dominate and dictate the overall execution time. Synchronous methods from previous works apply the same scheduling model, although often without explicitly referring to BSP, XPRAM, or supersteps.
3. Barrier list schedulers
The idea of adapting list schedulers to a setting with communication barriers has been explored recently by Papp et al. [PAKY24] in the context of abstract BSP scheduling. The barrier-list-scheduler prototype in their paper behaves similarly to a classical list scheduler, assigning concrete time steps to the starting and finishing times of each task. Whenever a core $p$ becomes free, it prioritizes assigning a next vertex to this core from those that are only computable on $p$ in the current superstep, since some of its parents were computed on $p$ in the current superstep. When half of the cores become idle (i.e., cannot be assigned new vertices without communication), a synchronization barrier is inserted.
The skeleton of our schedulers is similar to the base idea of this algorithm, but with many of the ingredients improved. In particular, the essence of the algorithm lies in the priority function that selects the next vertex to assign to a free core when there are multiple vertices ready. We present two schedulers with priority functions tailored towards the problem of developing work-balanced schedules with few synchronization steps. Both of our schedulers define this priority function via assigning priority scores to every vertex that is computable in the current superstep. This score may be core-specific as a vertex may be computable on more than one core. Whenever a core becomes free, our algorithms select the vertex with the highest score for the corresponding core as the next vertex to compute on this core.
Data: A vertex-weighted DAG $G=(V,E,\omega)$, a set of cores $P$, and a parameter $\alpha\in\>]0,1]$.

### v2

@@ Abstract.
We develop and analyze new scheduling algorithms for solving sparse triangular linear systems (SpTRSV) in parallel. Our approach produces highly efficient synchronous schedules for the forward- and backward-substitution algorithm. Compared to state-of-the-art baselines HDagg [ZCL+22] and SpMP [PSSD14], we achieve a 3.32×3.32$\times$3.32 × and 1.42×1.42$\times$1.42 × geometric-mean speed-up, respectively. We achieve this by obtaining an up to 12.07×12.07$\times$12.07 × geometric-mean reduction in the number of synchronization barriers over HDagg, whilst maintaining a balanced workload, and by applying a matrix reordering step for locality. We show that our improvements are consistent across a variety of input matrices and hardware architectures.
@@
22footnotetext: Joint first authors; listed in alphabetical order.
@@ 1. Introduction
Systems of linear equations are ubiquitous and solving them fast numerically with high accuracy is essential to engineering, big data analytics, artificial intelligence, and various scientific fields. Key techniques in scaling to ever larger linear systems have been exploiting the sparsity of non-zero coefficients in modern algorithms, as well as leveraging the multi-core or multi-processor architectures of high-performance computing systems. However, whilst sparsity reduces computational load, the typically irregular distribution of non-zero elements complicates the development of efficient parallel algorithms, as the lack of structure hinders workload balancing and limits the ability to minimize communication between processors.
In this paper, we concern ourselves with solving sparse triangular systems of linear equations (SpTRSV) using parallel machines; i.e., solving a linear system $Lx=b$, where $L$ is a sparse triangular matrix and $b$ is a dense vector. Although solving sparse triangular linear systems marks a special case, it often arises as an important step in procedures solving more general linear systems. Some concrete examples are (sparse) LU, QR, and Cholesky decompositions, Gauß–Seidel, and so forth. Efficient parallel-computation schedules for SpTRSV are of particular importance in applications where the same sparsity pattern is used repeatedly. Such is the case in simulations of various physical systems, for instance, ones that are based on the finite element method on a fixed mesh.
@@
Satisfying both of these needs simultaneously has proven to be challenging due to the irregular interdependence of computed values and the fine-grained nature of the problem. Early algorithms include so-called wavefront schedulers [AS89, Sal90], which repeatedly schedule all computations whose prerequisites are met, known as the wavefronts, cf. Figure 1(b), followed by a synchronization barrier. They, however, suffer from large overhead stemming from frequent global synchronization [PSSD14]. Similarly, early asynchronous approaches such as self-scheduling [SMB88] had the drawback of incurring overheads due to numerous fine-grained synchronizations [RG92].
@@ 1.1. Our contribution
Our work continues along the same path of reducing the number of synchronization barriers. We present and analyze a new scheduling algorithm named GrowLocal which is tailored specifically towards the SpTRSV application. In our experiments, we establish that this algorithm produces significantly superior parallel schedules compared to the baseline methods. Specifically, GrowLocal achieves a reduction in execution time of 1.42×1.42$\times$1.42 × compared to SpMP and of 3.32×3.32$\times$3.32 × compared to HDagg, on the SuiteSparse Matrix Collection benchmark [DH11] used by previous studies, see Figure 1.2. We further evaluate GrowLocal on pre-processed variants of the SuiteSparse matrices motivated by applications, and observe speedups of up to 1.80×1.80$\times$1.80 × and 2.20×2.20$\times$2.20 × over SpMP and HDagg, respectively. On synthetic random matrices that are hard to schedule by design, the difference to the baselines is even larger: the algorithm achieves a speed-up of 2.50×2.50$\times$2.50 × compared to SpMP and 10.12×10.12$\times$10.12 × compared to HDagg in execution time.
Figure 1.2. Geometric mean and interquartile ranges of speed-ups over Serial of our algorithms on the SuiteSparse Matrix Collection [DH11] on an Intel x86 machine using 22 cores.
The algorithm obtains these speed-ups by significantly reducing the number of synchronization barriers required: we report a 12.07×12.07$\times$12.07 × reduction in the number of barriers relative to HDagg on the SuiteSparse data set, whilst maintaining a good workload balance. The results also show that our scheduler provides consistent improvements over several different computing architectures and types of input matrices. The running time of the scheduling algorithm itself is also comparable to the state-of-the-art baselines, making it a viable tool for various applications.
@@
  • a novel algorithm for generating efficient parallel schedules for SpTRSV execution;
@@
  • extensions of previous DAG coarsening techniques that enhance the schedules, and a short theoretical proof that these preserve acyclicity; and
@@
  • experiments confirming that the above schedulers achieve significant speed-ups over the SpMP and HDagg baselines, on various architectures and data sets, including an ablation study of the individual techniques proposed.
1.1.1. GrowLocal scheduling algorithm
There are numerous prior works on parallel DAG scheduling in the literature. When the number of cores is limited, the best results are often achieved by so-called list scheduling algorithms [Gra69, ACD74, HCAL89, RVG02, MSQ03], which schedule the vertices in a topological order according to some priority function [WS18]. On the other hand, DAG scheduling with barrier synchronization is a somewhat different setting, and there are only a few previous works that address this problem. One state-of-the-art example here is the HDagg algorithm mentioned before [ZCL+22], which can also be interpreted as scheduler for general DAGs. Besides this, the idea of adapting list schedulers to a barrier synchronization setting has also been explored recently by Papp et al. [PAKY24] for abstract bulk-synchronous-parallel (BSP) scheduling.
Our GrowLocal algorithm takes a rather different approach than these previous methods, but also incorporates some of their underlying strengths. On a high level, GrowLocal considers a parameter $\alpha$, and tries to form the part of the schedule until the next synchronization barrier (the next so-called superstep) by assigning approximately $\alpha$ vertices to each of the cores before this next barrier. The parameter $\alpha$ is then iteratively increased, examining larger parts of the DAG as the potential next superstep, as long as this is possible while also ensuring a sufficient amount of parallelization between the cores.
During the development of the schedule, the algorithm always maintains the set of vertices that are ready to be executed, i.e., all their parents have been computed. At any point during the algorithm, if we consider the current superstep, such a ready vertex $v$ may be executable either on any of the cores (if all parents of $v$ were computed before the last barrier), only on a specific core $p$ (if a parent of $v$ was computed on $p$ since the last barrier), or on none of the cores (if we computed parents of $v$ on multiple cores since the last barrier). When selecting the next vertices to assign to a core $p$ during our algorithm, GrowLocal first prioritizes vertices that can only be executed on the core $p$ before the next barrier. This is inspired by the scheduler of [PAKY24], and it ensures that we can compute significantly more vertices before having to insert a new barrier.
Apart from this, GrowLocal simply selects vertices to assign to a core based on their IDs in order to group neighboring vertices onto the same core and superstep. This leads to significantly better locality for the developed schedule than in case of, e.g., list schedulers, and this has a large positive impact on the overall performance of the SpTRSV computation.
@@ 1.1.2. Coarsening
Graph coarsening techniques are widely applied in graph partitioning tools [IKS75, KAKS97, Sch20], where they greatly reduce the size of the graph and improve data locality. These techniques can also be applied to DAG scheduling [PSSS21, ZCL+22], where they can further help to reduce the number of synchronization steps on top of the aforementioned benefits. Following the coarsening, the scheduling algorithm is applied to the coarse graph and the resulting schedule is subsequently pulled back to the original graph to obtain the final schedule. In order to produce a valid scheduling problem, the coarsening methods are required to preserve the acyclicity of DAGs. Methods that fulfill this criteria have been studied in several works before, see, for example, [CLB94, FER+13, HKU+17, ZCL+22] and references therein.
In Section 4, we introduce the concept of cascades to generalize the coarsening techniques utilized in [CLB94, §4] and [ZCL+22, §IV.B]. We then formally prove that coarsening techniques based on cascades always preserve acyclicity. In Section 7.3, we evaluate the effect of the coarsening algorithm developed in Section 4 on our scheduling algorithm, GrowLocal.
@@ 1.1.3. Reordering
Besides the algorithms above, we also apply a matrix reordering step to drastically improve data locality during the SpTRSV computation. Specifically, once the schedule is developed, we symmetrically permute the matrix according to the schedule, ensuring that values computed after each other on the same core are close to each other in this permuted representation. This idea has already been explored by Rothberg–Gupta in the 1990s [RG92], but it has not been applied in modern SpTRSV baselines, which instead try to make use of existing data locality when deriving a schedule.
1.1.4. Block parallel scheduling
A known optimization technique for parallel SpTRSV execution is to break the lower triangular matrix into blocks [AS89, May09, AYU21, YSAU20]. These blocks may be on the diagonal, which corresponds to a smaller instance of (sparse) triangular solve, or completely off the diagonal, which corresponds to a (sparse) matrix-vector multiplication. The separation of the easily parallelizable (sparse) matrix-vector-multiplication blocks from the hard to parallize (sparse) triangular blocks has been particularly impactful for GPU implementations [LLH+16, LNL20].
In this paper, we use this block decomposition to run the GrowLocal scheduling algorithm in parallel on each (sparse) triangular block. The resulting synchronous schedules can then be combined one after the other (with a synchronization barrier between individual schedules) to a schedule for the whole triangular matrix. This leads to a super-linear speed-up in the scheduling time whilst having a moderate effect on the parallel SpTRSV solve time.
@@ 1.2. Additional related work
Besides the forward-/backward-substitution algorithm, there are also other methods for solving sparse triangular systems, for example inversion. For this method, we mention the memory-optimal algorithms developed in previous works [AS93, PA92].
@@ 2.1. Graph notation
We model our computations as a directed acyclic graph (DAG) $G=(V,E)$, which consists of a set of vertices $V$ and a set of directed edges $E\subseteq V\times V$. For any vertex $v\in V$, the sets of vertices $\{u\,|\,(u,v)\in E\}$ and $\{u\,|\,(v,u)\in E\}$ are called the parents of $v$ and the children of $v$, respectively. The in- and out-degree of $v$, denoted by ${\rm deg}^{-}(v)$ and ${\rm deg}^{+}(v)$, respectively, are the number of parents and children of $v$. The degree of $v$, denoted by ${\rm deg}(v)$, is the sum of its in- and out-degree. If a vertex of the DAG has no parents/children, then it is called a source/sink vertex, respectively. The DAG in our model is also complemented by vertex weights $\omega:V\rightarrow\mathbb{Z}_{>0}$ to indicate the compute cost of each operation.
@@ 2.2. Problem definition and notation
When solving sparse triangular systems, we are given a triangular matrix $A=(A_{i,j})_{i,j=1,\dots,n}\in\mathbb{R}^{n\times n}$, a dense vector $b=(b_{1},\dots,b_{n})^{T}\in\mathbb{R}^{n}$, and the goal is to solve the equation $Ax=b$ for the vector $x=(x_{1},\dots,x_{n})^{T}\in\mathbb{R}^{n}$. We assume that $A$ is non-singular, such that all its diagonal elements are non-zero. In case of a lower triangular matrix $A$, there is a natural forward-substitution algorithm for the problem, which iterates through the rows of $A$ in order and computes the values of $x$ as $x_{1}=\tfrac{b_{1}}{A_{1,1}}$, $x_{2}=\tfrac{b_{2}-A_{2,1}x_{1}}{A_{2,2}}$, and, in general, as
@@
In case of an upper triangular matrix $A$, a backward-substitution algorithm follows symmetrically in the reverse direction.
@@
The parallel execution of this DAG then directly corresponds to a parallel execution of the SpTRSV. Many previous works found it more convenient to discuss their scheduling methods for this problem using this DAG representation.
@@
The total cost of a schedule is determined by the workload balance within each superstep and the number of synchronization barriers. The original BSP model includes also communication volume in its cost function. For the SpTRSV application, however, the communication happens in parallel to the computation and resolving the synchronizations. Hence, the latter two dominate the overall execution time. Synchronous methods from previous works apply the same scheduling model, although often without explicitly referring to BSP, XPRAM, or supersteps.
3. The GrowLocal scheduler
Our GrowLocal algorithm is tailored specifically to the DAG scheduling problem with synchronization barriers. The algorithm forms the supersteps one by one, always aiming to make the current superstep as large as possible while maintaining a good workload balance. The current superstep is formed through several iterations with a superstep length parameter $\alpha$. The algorithm attempts to form a new superstep with approximately $\alpha$ vertices assigned to each core, and gradually increases $\alpha$ as long as this allows sufficient parallelization.
Specifically, in a single iteration with parameter $\alpha$, the algorithm first assigns (up to) $\alpha$ vertices to the first core, and considers the sum $\Omega_{1}$ of the weights of these vertices. It then assigns vertices up to total weight of at most $\Omega_{1}$ to the second core, third core, and so forth. Let $\Omega_{p}$ denote the total weight allocated to core $p$ in this iteration. We associate a parallelization score of
$\beta=\frac{\sum_{p}\Omega_{p}}{\max_{p}\Omega_{p}+L}$ (3.1)
to the current iteration. Here, $L$ is a parameter reflecting the penalty (time cost) incurred by each new synchronization barrier [^1] The value of $L$ may be architecture dependent. In this study, we set $L=500$ based on synchronization cycles and a small empirical evaluation.. In order to consider the superstep allocation of the current iteration worthy, the algorithm requires that its score $\beta$ is relatively large, i.e., close to the parallelization score achieved in the previous iterations.
In order to form a superstep, the algorithm begins with a minimal length $\alpha=20$ iteration. This first iteration is always considered worthy, regardless of its parallelization score. Then, in each subsequent iteration, we consider a different choice for the next superstep: the assignments of the previous iteration are invalidated, the parameter $\alpha$ is increased by a factor of $1.5$, and a new potential superstep is formed, with more vertices assigned to each core. If the resulting parallelization score is still high enough, then the superstep allocation of this iteration is also considered worthy, and the process continues. Otherwise, the last worthy superstep allocation is finalized as the current superstep. The high-level pseudocode of the algorithm is outlined in Algorithm 3.1.
Naturally, when assigning vertices to a specific core $p$ in a superstep, there may be numerous ready-to-compute vertices that we can choose from, and selecting among these is a key aspect to any scheduler. Similarly to the heuristic of [PAKY24], our algorithm first prioritizes those vertices that are only computable on $p$ in this superstep, since some of their parents were assigned to $p$ in the current superstep. In lack of such vertices, GrowLocal simply selects the vertices with smallest IDs.
We note that while this ID-based selection may seem simple at first, it in fact plays a crucial role in the success of our algorithm. Previous scheduling heuristics usually assign vertices to the different cores simultaneously in order to ensure work balance. In contrast to this, GrowLocal first assigns vertices to the first core, then to the second core, and so forth. With the ID-based selection, this often leads to schedules where the vertices on a core are more-or-less consecutive blocks in the matrix, which drastically improves locality during the computation. This is especially important in matrices from applications, which are often already ordered superbly with respect to locality, and thus preserving this is crucial. As such, GrowLocal can also be loosely understood as a method combining the strengths of previous DAG schedulers with barrier synchronization: similarly to [PAKY24], it allows a priority-based choice between vertices when forming a superstep, but as in [ZCL+22], it preserves locality by aiming to assign consecutive vertices to the same core.
Data: A vertex-weighted DAG $G=(V,E,\omega)$ and a set of cores $P=\{1,2,\dots,k\}$.